The reaction N2(g)+O2(g)⇌2NO(g)  is carried out at a temperature at which Kc = 0.050. The reaction mixture starts with only the product, [NO] = 0.0300 M , and no reactants.

Part A

Find the equilibrium concentrations of N2 at equilibrium.

Express your answer to two significant figures and include the appropriate units.

[N2] =

Part B

Find the equilibrium concentrations of O2 at equilibrium.

Part C

Find the equilibrium concentrations of NO at equilibrium.

Express your answer to two significant figures and include the appropriate units.
Answer #1

Part A

Find the equilibrium concentrations of N2 at equilibrium.

Express your answer to two significant figures and include the appropriate units.

The reaction N2(g)+O2(g)⇌2NO(g)  

Kc = [NO]^2/[N2][O2]= 0.050

Now ICE table

               N2(g)+O2(g)⇌2NO(g)  

I             0               0     0.0300                 

C          +x/2            +x/2 -x

E           x/2           x/2       0.0300-x

Now kc,

[NO]^2/[N2][O2]= 0.050

(0.0300-x)^2/ x/2*x/2= 0.050

X = 0.027 M

[N2] = X/2= 0.027/2=0.0135M

Part B

Find the equilibrium concentrations of O2 at equilibrium.

[O2] = X/2= 0.027/2=0.0135M

Part C

Find the equilibrium concentrations of NO at equilibrium.

Express your answer to two significant figures and include the appropriate units.

[NO]= 0.0300-x = 0.0300-0.027=0.003M

Cross check

Now kc,

[NO]^2/[N2][O2]= 0.050

= (0.003)^2/0.0135*0.0135

=0.050